# Safely Learning with Private Data: A Federated Learning Framework for Large Language Model

### v2

@@ Abstract
Private data, being larger and quality-higher than public data, can greatly improve large language models (LLM). However, due to privacy concerns, this data is often dispersed in multiple silos, making its secure utilization for LLM training a challenge. Federated learning (FL) is an ideal solution for training models with distributed private data, but traditional frameworks like FedAvg are unsuitable for LLM due to their high computational demands on clients. An alternative, split learning, offloads most training parameters to the server while training embedding and output layers locally, making it more suitable for LLM. Nonetheless, it faces significant challenges in security and efficiency. Firstly, the gradients of embeddings are prone to attacks, leading to potential reverse engineering of private data. Furthermore, the server’s limitation of handle only one client’s training request at a time hinders parallel training, severely impacting training efficiency. In this paper, we propose a Federated Learning framework for LLM, named FL-GLM, which prevents data leakage caused by both server-side and peer-client attacks while improving training efficiency. Specifically, we first place the input block and output block on local client to prevent embedding gradient attacks from server. Secondly, we employ key-encryption during client-server communication to prevent reverse engineering attacks from peer-clients. Lastly, we employ optimization methods like client-batching or server-hierarchical, adopting different acceleration methods based on the actual computational capabilities of the server. Experimental results on NLU and generation tasks demonstrate that FL-GLM achieves comparable metrics to centralized chatGLM model, validating the effectiveness of our federated learning framework. [^1] The code is avaliable at: https://github.com/TAP-LLM/SplitFedLLM
@@ 1 Introduction
Experimental results on NLU and generation tasks demonstrate that FL-GLM achieves performance comparable to centralized chatGLM-6B models, validating the effectiveness of our framework. Further analysis of training costs indicates that our client-batch and server-hierarchical mechanisms can save more than 48% of training time.
@@ 2.2 Attacks and Defenses
In federated learning, various eavesdroppers threaten client privacy, including servers attempting data recovery and peer-clients intercepting data sent to servers.In NLP, attacks from embedding gradients can easily recover users’s private data.  (Gupta et al., 2022) proposes to infer which words the client used by observing the non-zero values in embedding gradients. They then use beam search and resort to arrange these words, thereby reconstructing private data.
To counter this, they recommend freezing embedding layers during training.  (Zhu et al., 2019) briefly mentions defending by adding differentially private noise or setting small gradients to zero (gradient clipping).  (Huang et al., 2020) propose MixUp data augmentation on the BERT model’s [CLS] token.  (Yaldiz et al., 2023) suggest server-side cosine similarity checks on client-uploaded weights to filter out malicious clients. However, these defenses often reduce model accuracy (Yu et al., 2021; Li et al., 2021).
@@ 3.1 Model Split
For protecting data privacy, the FL-GLM framework splits LLM into three parts for deployment. During forward operations, the client-side model processes private data to generate smashed data, which is then sent to the server-side model for computation. Encrypting the smashed data ensures its security. Given the input data $x=\left\{x_{1},\dots,x_{L}\right\}$ and the next output $y$, the smashed data $h_{0}$ of the client is defined as:
@@ 3.2 Encrypted Transmission
3) Compute the Euler function $\phi\left(n\right)=\left(p-1\right)\left(q-1\right)$.
@@ 4.3.3 Security Analysis
The experimental results are shown in Table 4. When the client only has the embedding layer like FedBert model, $F^{-1}$ is a single Transformer block, the attack model can achieve a BLEU-4 score of 28.570 and a ROUGE-1 score of 33.290, while in the FL-GLM framework, where the client contains the embedding layer and an LLM-Block, $F^{-1}$ is a single layer Transfomer, all the metrics of the attack model are all close to 0. Therefore, the security of FL-GLM could be proven in experiments. Additionally, we find that the attack metrics’ performance of a single-block Transformer is similar to that of a multi-block Transformer. Therefore, the optimal split point, based on experimental results, might be a single-block Transformer, even though it is challenging to prove theoretically.
@@ Appendix C Impact of Average Period
The results are shown in Table  6, where the model with an average period of 100 steps slightly outperforms the model with an average period of 50 steps in the BoolQ task. However, in the WiC, RTE, and MultiRC tasks, better results are achieved with an average period of 50 steps. In the COPA and CB tasks, the averaging period has no effect on performance. The most noticeable difference occurs in the WSC task, with scores of 71.2 and 66.3 for an average period of 50 steps and 100 steps, respectively, for serial training, 63.5 and 65.4 for client-batch parallel, and flat accuracy scores for server-hierarchical. Among all the evaluation tasks, the WSC task has the highest sensitivity to the average period, but the average training period has little effect on the overall performance of the FL-GLM model with the same training strategy.

### v3

@@ Abstract
Private data, being larger and quality-higher than public data, can greatly improve large language models (LLM). However, due to privacy concerns, this data is often dispersed in multiple silos, making its secure utilization for LLM training a challenge. Federated learning (FL) is an ideal solution for training models with distributed private data, but traditional frameworks like FedAvg are unsuitable for LLM due to their high computational demands on clients. An alternative, split learning, offloads most training parameters to the server while training embedding and output layers locally, making it more suitable for LLM. Nonetheless, it faces significant challenges in security and efficiency. Firstly, the gradients of embeddings are prone to attacks, leading to potential reverse engineering of private data. Furthermore, the server’s limitation of handle only one client’s training request at a time hinders parallel training, severely impacting training efficiency. In this paper, we propose a Federated Learning framework for LLM, named FL-GLM, which prevents data leakage caused by both server-side and peer-client attacks while improving training efficiency. Specifically, we first place the input block and output block on local client to prevent embedding gradient attacks from server. Secondly, we employ key-encryption during client-server communication to prevent reverse engineering attacks from peer-clients. Lastly, we employ optimization methods like client-batching or server-hierarchical, adopting different acceleration methods based on the actual computational capabilities of the server. Experimental results on NLU and generation tasks demonstrate that FL-GLM achieves comparable metrics to centralized chatGLM model, validating the effectiveness of our federated learning framework.
@@ 1 Introduction
Experimental results on NLU and generation tasks demonstrate that FL-GLM achieves performance comparable to centralized chatGLM-6B models, validating the effectiveness of our framework. Further analysis of training costs indicates that our client-batch and server-hierarchical mechanisms can save more than 48% of training time. Our code is available at: https://github.com/TAP-LLM/SplitFedLLM.
@@ 2.2 Attacks and Defenses
In federated learning, various eavesdroppers threaten client privacy, including servers attempting data recovery and peer-clients intercepting data sent to servers.In NLP, attacks from embedding gradients can easily recover users’s private data.  (Gupta et al., 2022) proposes to infer which words the client used by observing the non-zero values in embedding gradients. They then use beam search and resort to arrange these words, thereby reconstructing private data. To counter this, they recommend freezing embedding layers during training.  (Zhu et al., 2019) briefly mentions defending by adding differentially private noise or setting small gradients to zero (gradient clipping).  (Huang et al., 2020) propose MixUp data augmentation on the BERT model’s [CLS] token.  (Yaldiz et al., 2023) suggest server-side cosine similarity checks on client-uploaded weights to filter out malicious clients. However, these defenses often reduce model accuracy (Yu et al., 2021; Li et al., 2021).
@@ 3.1 Model Split
For protecting data privacy, the FL-GLM framework splits LLM into three parts for deployment. Take ChatGLM as an example, the complete ChatGLM model contains the embedding layer, 28 ChatGLM blocks and the final linear layer, the client side contains the embedding layer, the 0th ChatGLM, the 27th ChatGLM block and the final linear layer, and the 1st to 26th ChatGLM are deployed on the server side, each block is a transformer structure. During forward operations, the client-side model processes private data to generate smashed data, which is then sent to the server-side model for computation. Encrypting the smashed data ensures its security. Given the input data $x=\left\{x_{1},\dots,x_{L}\right\}$ and the next output $y$, the smashed data $h_{0}$ of the client is defined as:
@@ 3.2 Encrypted Transmission
3) Compute the Euler’s totient function $\phi\left(n\right)=\left(p-1\right)\left(q-1\right)$. For a prime number $p$, there are $p-1$ numbers that are coprime with $p$; similarly, for a prime number $q$, there are $q-1$ numbers that are coprime with $q$. Since $p$ and $q$ are coprime with each other, the Euler’s totient values for $p$ and $q$ can be multiplied directly.
@@ 4.3.1 Training Efficiency
Furthermore, we measured training communication time between two computers (server and client) within the same LAN with a bandwidth of 1100MB///s. Results indicate that training across two machines in the LAN is approximately five times slower than centralized training (centralized training: 0.91s///step, stimulate training FL-GLM in single machine with two GPUs: 1.79s///step, training FL-GLM with two machines: 4.83s///step).
@@ 4.3.3 Security Analysis
The experimental results are shown in Table 4. When the client only has the embedding layer like FedBert model, $F^{-1}$ is a single Linear layer, the attack model can achieve a BLEU-4 score of 28.570 and a ROUGE-1 score of 33.290, while in the FL-GLM framework, where the client contains the embedding layer and an LLM-Block, $F^{-1}$ is a single layer Transfomer, all the metrics of the attack model are all close to 0. Therefore, the security of FL-GLM could be proven in experiments. Additionally, we find that the attack metrics’ performance of a single-block Transformer is similar to that of a multi-block Transformer. Therefore, the optimal split point, based on experimental results, might be a single-block Transformer, even though it is challenging to prove theoretically.
Table 4: Security Analysis
4.3.4 More Basement Model
To further validate the practicality and generalizability of our FL-GLM framework, we incorporated Llama2-7B-Chat (Touvron et al., 2023) as a baseline model. We split Llama2-7B-Chat according to the FL-GLM framework, resulting in a variant denoted as FL-Llama. The performance evaluation experiment was conducted utilizing six datasets from the SuperGLUE benchmark with serial training strategy and utilizing the LoRA to fine-tune. Note that Llama2-7B-Chat supports instruction fine-tuning, leading us to adapt the input prompts for each dataset in accordance with the Meta’s official recommended instruction templates, rather than employing the Cloze question templates in Appendix B. The experimental results are presented in Table 5, which indicate that our proposed framework could achieve comparable accuracy metrics, and suggest that the FL-GLM framework is agnostic to the base model type and does not significantly affect performance.
Furthermore, we validated the FL-GLM framework on the Chinese medical dataset Huatuo-26M (Li et al., 2023), referring to the Llama2-Chinese-7B-Chat-based model as FL-Chinese-Llama. The full Huatuo-26M dataset contains 2,623,904 Q&A pairs in the training dataset and 264,041 Q&A pairs in the test dataset. We randomly sampled 3,000 Q&A pairs from the full dataset for training and 300 Q&A pairs from the test set for evaluation. It should be noted that, since the limited Chinese language capability of Llama2-7B-chat, we chose Llama2-Chinese-7B-Chat  [^3] https://huggingface.co/FlagAlpha/Llama2-Chinese-7B-Chat as the baseline model for this experiment, which has been fine-tuned on a large Chinese corpus. As illustrated in Table 6, the proposed training framework maintains model performance that is closely comparable to centralized training. Specifically, the performance metrics of FL-GLM exhibit a degradation of no more than 0.72 when compared to ChatGLM-6B, while FL-Chinese-Llama shows a decline of no more than 0.38 when compared to Llama2-Chinese-7B-Chat. Whereas, the inter-model comparisons show that the Chinese conversational capabilities of ChatGLM significantly surpass those of Llama2-Chinese-7B-Chat. Therefore, we recommend using ChatGLM as the base model for FL-GLM framework in Chinese application scenarios.
Model CB COPA RTE WiC BoolQ WSC Llama2-7B-Chat 85.71 75.00 73.29 70.85 83.79 68.27 FL-Llama(ours) 85.71 75.00 73.29 70.83 82.35 68.37
Table 5: FL-GLM’s performance on Llama2-7B-Chat
Model BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L ChatGLM-6B 9.32 29.94 8.89 24.07 FL-GLM(OURS) 8.95 29.22 8.61 23.40 Llama2-Chinese-7B-Chat 3.07 19.42 2.95 15.08 FL-Chinese-Llama(ours) 3.04 19.04 2.79 14.85
Table 6: Results on sampled Huatuo-26M test set.
@@ Appendix C Impact of Average Period
The results are shown in Table  8, where the model with an average period of 100 steps slightly outperforms the model with an average period of 50 steps in the BoolQ task. However, in the WiC, RTE, and MultiRC tasks, better results are achieved with an average period of 50 steps. In the COPA and CB tasks, the averaging period has no effect on performance. The most noticeable difference occurs in the WSC task, with scores of 71.2 and 66.3 for an average period of 50 steps and 100 steps, respectively, for serial training, 63.5 and 65.4 for client-batch parallel, and flat accuracy scores for server-hierarchical. Among all the evaluation tasks, the WSC task has the highest sensitivity to the average period, but the average training period has little effect on the overall performance of the FL-GLM model with the same training strategy.